OpenStudy (thatemokid66):

I have been working on this problem for hours and still can't get it. Can anyone help? It's not multiple choice so I'd need help working it out. Use the binomial expression (p + q)^n to calculate a binomial distribution with n = 5 and p = 0.3.

OpenStudy (thatemokid66):

This answer is not correct because you it is (p +q) ^n not (p+q) * 5

OpenStudy (bartyrocks3):

Alright. Give me a couple of minutes to try to work this out. brb

OpenStudy (thatemokid66):

no problem, thank you :)

OpenStudy (bartyrocks3):

(p + q)^5 = p^5 + (5p^4)(q) + (10p^3)(q^2) + (10 p^2)(q^3) + 5p(q^4) + q^5 If p = 0.3, then q = 1 - p = 1 - 0.3 = 0.7. I`ll let you finish the problem.

OpenStudy (bartyrocks3):

my other account got deleted due to inactivity. Now I am stuck with restarting all over again. I had a smart score of 79...

OpenStudy (thatemokid66):

Thank you so much :) and dang that sucks :( but considering how easily you helped me I think you'll get them back quickly no worries :3

OpenStudy (bartyrocks3):

no problem.